# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v5

@@ 9.2 Supplementary Notes and Extended Discussion
A well documented example of the effectiveness of mission command in facilitating adaptation is Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s. Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [ingesson2016].
S8. From analogy to evidence: candidate empirical and simulation tests
The main text uses analogies (mission command, economic delegation, and distributed swarms) as intuition-building thought experiments rather than as evidence. To turn the delegation hypothesis into a research programme, one can state testable predictions and compare systems under controlled variation of delegation depth.
@@
Construct an agent-based model of a collective where (i) individuals have local perception and local policy spaces, (ii) a collective objective is defined at a higher level, and (iii) a “delegation parameter” controls how much decision-making is centralised versus left to individuals (e.g. by constraining local policy spaces or routing decisions through a leader). Expose the collective to non-stationary environments (distribution shifts, component failures) and measure: sample efficiency (data required to regain acceptable performance), robustness (graceful degradation under failures), and identity preservation (whether the collective policy remains coherent rather than splintering into incompatible sub-collectives). Theorem 5.1 predicts that as delegation is pushed upward (lower layers become static or over-constrained), achievable utility at higher layers should become bounded by the reduced weakness of the lower-layer realised policies.
@@
Compare learning systems that differ only in where adaptation is permitted. For example, keep a fixed high-level learning rule but vary whether lower layers are static (fixed hardware / fixed microcode) versus adaptive (e.g. reconfigurable hardware, adaptive firmware, or reconfigurable substrates). Measure sample and energy efficiency on tasks with distribution shift and compositional generalisation (e.g. ARC-style tasks). This isolates the “where in the stack” hypothesis from the choice of algorithm at the top.
@@
For hybrid agents such as simulated systems of human and AI agents, where the goal is to “align” the AI agents with goal goals, systematically vary the boundary conditions (stimulation regimes, feedback mappings, resource constraints) and quantify the trade-off between task performance and preservation of collective identity (non-empty intersections of viable policy sets, as in the organoid formalisation above). The design expectation is that weaker (less exclusionary) boundary conditions should preserve a larger viable policy space while still enabling the desired interface-level behaviour.
@@
Implement Active Inference-style control at one or more layers, but run the same comparison as above: (i) active inference on a fixed substrate versus (ii) active inference on a substrate that permits deeper delegation (adaptation in lower layers). This helps disentangle improvements due to the inferential objective from improvements due to stack-level delegation.
These designs are intended as starting points rather than definitive protocols; the unifying aim is to replace analogy with controlled variation of delegation, and to measure adaptability in the efficiency-of-adaptation sense used throughout this paper.

### v6

@@ Acknowledgements
A well documented example of the effectiveness of mission command in facilitating adaptation is Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s. Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [47, 48].
Strategic-level intent and a small set of constraints are communicated downward, while tactical-level decision-makers fill in the remaining degrees of freedom on the ground. In this reading, the bridge between leadership and soldiers is not a detailed plan but a constraint system that keeps local improvisation aligned with higher-level purpose. The purpose of this note is not to treat Nordbat 2 as empirical evidence for Theorem 5.1, but to provide an intuition-building picture of what Stack Theory means by delegation [47, 48].
Mapped to the formalism, the strategic layer specifies a parent task $\omega$ (intent, mandate, and coordination requirements), while the tactical layer faces a family of child tasks $\alpha\sqsubset\omega$ that vary with local information and contingencies. Mission command enlarges the set of admissible completions by avoiding over-specification of the realised policy. It preserves a larger extension $|\mathrm{Ext}(\pi_{i})|$ at the lower layer, enabling more distinct higher-layer behaviours to be expressed one layer up (Theorem 5.1).
For hybrid and cyberphysical design, the analogy is that effective “control” can often be achieved by engineering boundary conditions (interfaces, reward signals, resource limits, safety constraints) rather than by prescribing micro-actions. Over-prescriptive constraints risk collapsing the intersection of viable component policies (loss of collective identity), whereas intent-plus-boundaries can preserve slack for local adaptation while maintaining coherence.
S8. Candidate empirical and simulation tests
The main text uses analogies (mission command, economic delegation, and distributed swarms) as intuition-building thought experiments rather than as evidence. It is primarily conceptual. This note lists candidate tests that target the core claim of Theorem 5.1. The goal is to compare systems under controlled variation of where adaptation is permitted in the stack.
@@
Construct an agent-based model of a collective. Include a delegation parameter that controls whether decisions are taken locally or routed through higher-level coordination. Expose the collective to abrupt environmental shifts. Measure sample efficiency, energy use, robustness, and identity preservation. Test whether deeper delegation improves efficiency in the sense predicted by reduced weakness of lower-layer realised policies.
@@
Compare learning systems that differ only in where adaptation is permitted. Hold the top-layer objective fixed and vary whether lower layers are static or reconfigurable. For example, compare fixed-substrate deployment against deployment that allows adaptation of firmware, routing, or morphology. Measure adaptation speed and compute the effective induced vocabulary size at each layer when possible. This isolates the stack-level delegation hypothesis from algorithm choice at the top.
@@
For hybrid agents such as simulated mixtures of human and AI agents or engineered organoids, treat alignment as hybrid agent design. Sweep boundary conditions that prune low-level policy spaces. Quantify the trade-off between task performance at the interface and preservation of non-empty collective policy intersection. Look for phase transitions where the intersection collapses, which corresponds to loss of collective identity.
@@
Implement Active Inference style control at one or more layers. Run the same comparison in an architecture that differs only in whether lower layers can adapt. This disentangles gains due to the inferential objective from gains due to deeper delegation.
S9. Extended applications and speculative directions
This note collects material that is peripheral to the main theorem but may be useful for future-work framing. These applications are intended as design hypotheses rather than as evidence for Theorem 5.1.
Organisational and economic systems.
Free-market trade is a great example of how delegation can facilitate adaptability in a large collective. As Hayek pointed out [44], not all of the information is available to a central planner and so control is best delegated to the individual. However, the accumulation of resources within my formalism is implicit, rather than explicit. I have been concerned only with whether a system satisfies whatever arbitrary notion of correct I choose (i.e. does it achieve homeostatic and reproductive goals in whatever amount of time I care about?), rather than how it evolves over time as a dynamical system. In other words, I have concerned myself only with whether a system arrives on time at the destination I want, not the journey it took to get there. Future research might extend Stack Theory to formalise time-series, or the accumulation of various resources. For example, a stack could represent a sequence of tasks over time where higher layers are later in time. Earlier I mentioned that task correctness can denote more holistic notions of emergent features robust to lower-level variation. A time-series model could describe the conditions under which such invariants emerge. It could be used to explore how the delegation of control, distribution of work and the accrual of resources change over time. For example under some conditions a free market may degrade into a maladaptive, top-down system. Those parts deprived of control or resources may then not be able to construct a policy that aligns with the larger whole, and so splinter off. To preserve the integrity of the system, policies which might support the desired invariants could be modelled and evaluated. Policy could be located along axes of distribution and delegation, like an alternative political compass. In a simple zero-sum game like a race, one example of such a policy is adaptive rubber banding [61]. A game mechanic (e.g. in Mario Kart) where success accrues risk as well as reward, keeping the game competitive.
Regulation and over-constraint risks.
The cancer analogy suggests a general caution. Guardrails that over-constrain a system may precipitate the failure mode they aim to prevent by collapsing the intersection of viable policies. For example, survey of 3,306 Germans found that restrictions promoting carbon-neutral lifestyles would trigger strong negative responses because they ‘restrict freedom’ [77]. This may be most relevant to proposals for safeguarded AGI and related governance approaches [29, 28, 3, 42], given the formal results pertain specifically to a generalisation-optimal model of intelligence [18]. These remarks are speculative and are intended as design hypotheses rather than claims about current regulatory regimes.